OpenStudy (anonymous):

Given normal random variable Z , with mean μ=22.8 and standard deviation σ=6.4 , we obtain a sample of Z of size 70 . Consider the new random variable Xˉ which is the average of our sample. 1) find the interval [μXˉ−a,μXˉ+a] symmetric about μXˉ satisfying P(Xˉ∈[μXˉ−a,μXˉ+a])=0.9 is given by

OpenStudy (anonymous):

those x are support to be \[\bar x\]\[\mu_{\bar x }\]

OpenStudy (tkhunny):

What Z-values puts (1-0.9)/2 = 0.10/2 = 0.05 in each of the symmetric tails?

OpenStudy (anonymous):

what do you mean ?

OpenStudy (anonymous):

you mean creating a graph ?